OpenStudy (anonymous):

Alright i did what i was supposed to when i factored this: 0 = x^2 – 10x – 24 and i got this when i did one of the first steps: (x^2 – 4x) – (6x - 24) How am i supposed to get this answer (x+12)(x-4) out of that anyone clarify my jumbled thoughts? :P lol

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

When you foil your result out you should get the original equation. A simpler equation that can be factored is x^2-25 = (x-5)(x+5), notice you get x^2-5x+5x-25, the 5x's cancel and you get the original equation. So when you're trying to get the factored version, logically you can do the simple arithmetic in your head. What two numbers when multiplied together give you 24 but equal -10 when foiled? well we have 2*12, which equals 24, and 2-12 = -10, so we can try subbing that in as : (x-12)(x+2)

OpenStudy (anonymous):

OHHH!!!! i instead of getting 2 and - 12 i got -4, and -6 because -4-6=-10.... WOW lol thanks so much!!!! :D

OpenStudy (anonymous):

no problem